OpenStudy (anonymous):

PLEASE HELP!!!!!!!! Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. (sin x) / (1-cos x) + (sin x) / (1+cos x) = 2 csc x

OpenStudy (saifoo.khan):

@kaiidayy What is this?

OpenStudy (saifoo.khan):

@Butterfly16 : \[\frac{\sin x}{1-\cos x} + \frac{\sin x}{1+\cos x}\]

OpenStudy (saifoo.khan):

Like this?

OpenStudy (anonymous):

Yes

OpenStudy (saifoo.khan):

\[\frac{\sin x(1+\cos x)+\sin x(1-\cos x)}{(1-\cos x)(1+\cos x)}\]

OpenStudy (saifoo.khan):

\[\frac{sinx +\sin x \cos x + \sin x - \cos x \sin x}{1 - \cos^2 x}\]

OpenStudy (saifoo.khan):

\[\frac{2 \sin x}{ 1 - \cos^2 x }\]

OpenStudy (saifoo.khan):

Can you solve now?

OpenStudy (anonymous):

Do they turn into a different identity?

OpenStudy (anonymous):

\[\frac{ 2 \sin x }{ 1-\cos^2 x }\]is equal to 2 csc x right?

OpenStudy (saifoo.khan):

Recall cos^2 x + sin^2 x =1

OpenStudy (anonymous):

I don't understand how to put that into the equation :/

OpenStudy (saifoo.khan):

cos^2 x + sin^2 x =1 sin^2x = 1 - cos^2 x

OpenStudy (saifoo.khan):

so: 2sinx/sin^2x

OpenStudy (anonymous):

And what would that go down to? I cant find an identity in my notes that has those..

OpenStudy (saifoo.khan):

this is the basic identity.

OpenStudy (anonymous):

Would I use sin(x)=1/csc(x)?

OpenStudy (saifoo.khan):

\[\Large \frac{2sinx}{\sin^2x } \to \frac{2 \sin x}{\sin x *\sin x}\]

OpenStudy (saifoo.khan):

cscx = 1/sinx Yes.

OpenStudy (anonymous):

so the sin x would cancel one on the top and give sin x?

OpenStudy (saifoo.khan):

you'll be left behind with: 2/sin x

OpenStudy (anonymous):

and that would be 2 csc x right? since 1/sin x is csc then 2/sin x would make it 2 csc x?

OpenStudy (saifoo.khan):

Yes sir.

OpenStudy (anonymous):

Thank you SO much for your help! I really truly appreciate it. :o)

OpenStudy (saifoo.khan):

:)